# CELA: Cost-Efficient Language Model Alignment for CTR Prediction

### v2

@@ Abstract.
Click-Through Rate (CTR) prediction holds a paramount position in recommender systems. The prevailing ID-based paradigm underperforms in cold-start scenarios due to the skewed distribution of feature frequency. Additionally, the utilization of a single modality fails to exploit the knowledge contained within textual features. Recent efforts have sought to mitigate these challenges by integrating Pre-trained Language Models (PLMs). They design hard prompts to structure raw features into text for each interaction and then apply PLMs for text processing. With external knowledge and reasoning capabilities, PLMs extract valuable information even in cases of sparse interactions. Nevertheless, compared to ID-based models, pure text modeling degrades the efficacy of collaborative filtering, as well as feature scalability and efficiency during both training and inference. To address these issues, we propose Cost-Efficient Language Model Alignment (CELA) for CTR prediction. CELA incorporates textual features and language models while preserving the collaborative filtering capabilities of ID-based models. This model-agnostic framework can be equipped with plug-and-play textual features, with item-level alignment enhancing the utilization of external information while maintaining training and inference efficiency. Through extensive offline experiments, CELA demonstrates superior performance compared to state-of-the-art methods. Furthermore, an online A/B test conducted on an industrial App recommender system showcases its practical effectiveness, solidifying the potential for real-world applications of CELA.
@@ 1. Introduction
The ID-based paradigm has garnered considerable success. Notable examples include DeepFM (Guo et al., 2017), PNN (Qu et al., 2016), and DCN (Wang et al., 2017), which focus on high-order explicit feature crossing, while DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) are geared towards mining behavior sequences. This paradigm, illustrated on the left side of Figure 1, entails encoding user and item features into sparse one-hot vectors, subsequently transformed into dense embedding vectors via a lookup process. Latent co-occurrences among features are discerned via a feature interaction layer. Over the past decade, the advancement of these models has centered on developing more sophisticated and scenario-adapted feature interaction layers as a means to enhance the overall performance. Despite their prevalence, ID-based models exhibit certain limitations: 1) Dependence on historical data: This dependence can result in suboptimal modeling in scenarios such as cold-start or long-tail elements, where historical interactions for items or features are limited or absent, leading to inadequate learning of embeddings. 2) Limited feature scope: Primarily focusing on categorical and numerical features, ID-based models overlook other modalities like text, which may harbor untapped knowledge that could potentially enhance the model’s effectiveness.
To alleviate these limitations, works such as PTab (Liu et al., 2022) and CTRL (Li et al., 2023) have integrated Pre-trained Language Models (PLMs) into recommender systems. These models employ hard prompts to structure raw features for each interaction, subsequently processed through PLMs for recommendation purposes, as shown on the right side of the Figure 1. By adopting natural language to represent features instead of one-hot vectors, this approach preserves the semantic integrity of features. This enables PLMs to apply their external knowledge and reasoning capabilities, effectively capturing valuable information even in cases of sparse interactions. For example, ID-based models struggle to learn embeddings for infrequently interacted items like bread, whereas PLMs utilize external knowledge to classify bread as a type of food, akin to toast. Concurrently, this natural language representation facilitates the integration of textual features, such as item descriptions, thereby augmenting the recommender system’s effectiveness.
However, these text-enhanced models face inherent limitations that hinder their practical deployment, also shown in Figure 1. First, modeling based exclusively on pure text may not be ideally suited for capturing collaborative signals, which often results in reduced performance (Li et al., 2023). Second, scalability issues emerge as the number of features increases, potentially leading to constructed text exceeding the language model’s maximum token limit and thus losing valuable information, particularly when the sequence of user behaviors is extensively long. Furthermore, it becomes impractical to delineate user behaviors in a detailed manner, including the specification of item features for each behavior. Last but not least, to bridge the gap between PLM outputs and recommendation tasks, PLMs are usually fine-tuned at the interaction-level (Liu et al., 2022; Geng et al., 2022; Li et al., 2023; Muhamed et al., 2021), yet this process is notably time-intensive. Such inefficiency inherent in PLM text processing extends to inference as well (Liu et al., 2022; Geng et al., 2022; Lin et al., 2023; Bao et al., 2023).
@@
Figure 1. ID-Based Models vs. Pre-trained Language Models (PLMs) for CTR prediction. The former converts tabular data into one-hot vectors, capturing collaborative signals without external knowledge. The latter populates data into a structured prompt, and then feeds the truncated text into a PLM, leveraging external world knowledge for data interpretation.
@@
To overcome these challenges, we propose the Cost-Efficient Language Model Alignment (CELA) for CTR prediction. CELA leverages textual features and PLMs to mitigate the cold-start issue and expand the feature scope. It integrates ID-based models to preserve the collaborative filtering effectiveness, and shifts away from the hard prompt mode, thereby addressing scalability concerns. Through item-level alignment, the PLM is fine-tuned effectively and oriented towards recommendation tasks. CELA is structured as a three-phase paradigm: 1) domain-adaptive pre-training stage, the PLM is further trained on a corpus of textual features, tailoring it to the specific dataset, while also training an ID-based model; 2) recommendation-oriented modal alignment stage, contrastive learning is employed to align the item text representations with the item feature embeddings of the ID-based model. This alignment is at the item level, ensuring minimal training overhead. Post-alignment, text representations of each item are cached, facilitating efficient access via item identifiers, thereby reducing inference costs; and 3) multi-modal feature fusion stage, these aligned text representations are integrated into a new ID-based model, along with non-textual features, enabling their interplay to identify potential click patterns. The final two stages are structured for alternate execution, ensuring ongoing alignment of the PLM with the progressively refined embedding table.
@@
  • We propose a novel paradigm for integrating textual features with ID-based models. This approach is model-agnostic and necessitates minimal modifications to the existing network architecture, rendering it conveniently applicable.
@@
  • Comprehensive offline experiments on public and industrial datasets demonstrate the effectiveness of CELA. An online A/B test within an App recommender system further validates CELA’s real-world efficacy and feasibility.
@@ 2.1. ID-Based Models for CTR Prediction
ID-based models have been predominant in industry recommendations over the past decade. They capture the characteristics of the items and users by one-hot vectors and learn the collaborative signals to predict the likelihood of a user clicking on an item (Lian et al., 2020). Their architectures are meticulously crafted, evolving into increasingly intricate ones. FM (Rendle, 2010) stands out as one of the earliest machine learning models to incorporate second-order feature interactions. Both WideDeep (Cheng et al., 2016) and DeepFM (Guo et al., 2017) incorporate deep components to learn high-level feature interactions. PNN (Qu et al., 2016) extends the feature interaction function and tightly couples the interaction of first-order and second-order features. AutoInt (Song et al., 2019) and DESTINE (Xu et al., 2021) incorporate self-attention mechanisms to weight different feature interactions. Fi-GNN (Li et al., 2019) leverages graph structures to model feature interactions. DCN (Wang et al., 2017), DCNv2 (Wang et al., 2021), and EDCN (Chen et al., 2021) capture meaningful feature crossings automatically, eliminating the need for manual feature engineering or exhaustive searching. DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) model user behavior sequences, aiming to capture the relationships between the target and historical behaviors.
@@ 2.2. Text-Enhanced Models for CTR Prediction
Language models have made remarkable strides, achieving unprecedented natural language understanding and generation. These models are pre-trained on massive datasets and adopt advanced architectures like Transformer (Vaswani et al., 2017), enabling diverse applications such as chatbots (Du et al., 2022), translation (Zhang et al., 2022), and content creation (Chung et al., 2022). Highlighting their potential in semantic comprehension, language models are considered promising for recommendation. Analyzing extra-textual elements such as ad text, search queries, or user reviews, these models may have the capability to discern user intent and subtle nuances between items, thereby enhancing ID-based ones. P5 (Geng et al., 2022) represents a paradigm that redefines various tasks into natural language processing (NLP) tasks. By structuring interactions and metadata in natural language and inputting them into T5 (Raffel et al., 2020), it can generate the target response. CTR-BERT (Muhamed et al., 2021) concatenates all features along with their respective feature names. It undergoes pre-training and distillation with BERT (Devlin et al., 2018). Subsequently, BERT encodes the text of users and items, generating representations that serve as inputs for a late fusion MLP. PTab (Liu et al., 2022) is pre-trained on BERT (Devlin et al., 2018) with a language head using the ”field: value” corpus, which is derived from tabular data. It is then fine-tuned for the CTR task, incorporating a classifier head. UniSRec (Hou et al., 2022) employs a frozen BERT (Devlin et al., 2018) as a text encoder and utilizes a Transformer (Vaswani et al., 2017) architecture to model user behaviors. CTRL (Li et al., 2023) uses contrastive learning for representation alignment between collaborative and semantic models at the interaction level. The collaborative model, now enriched with semantic information, is fine-tuned and can be deployed without the semantic model.
@@ 3.2. Embedding Layer
To compress those sparse, high-dimensional vectors into dense, lower-dimensional ones, the embedding layer is utilized. It acts as a lookup table, denoted by $\mathbf{E}^{(\cdot)}\in\mathbb{R}^{C^{(\cdot)}\times d}$ , which maps each categorical feature into a dense vector of a fixed size $d$. Here, $d$ represents the embedding size. Let $\boldsymbol{x}^{u}$ denote user profile, $\boldsymbol{x}^{h}$ denote browsing history, $\boldsymbol{x}^{c}$ denote contextual information, $\boldsymbol{x}^{i,n}$ denote non-textual features of target item $i$ and $\boldsymbol{x}^{h,n}$ denote non-textual features in user history. This process can be formalized as:
@@ 4. Methodology
  • Domain-Adapted Pre-training (DAP). It involves further pre-training of a PLM on a corpus comprising item-side text specific to the recommendation dataset, to better adapt the PLM to the unique characteristics of the dataset. Concurrently, a traditional ID-based model is trained.
@@
Notably, the last two stages are designed to be executed alternatively, where the PLM is continually aligned to the more advanced embedding table.
Figure 2. The overall framework of CELA. The first stage involves pre-training a PLM on domain-specific item text alongside training an ID-based model. In the second stage, item text representations from the PLM are aligned with the item-side feature embeddings of the ID-based model in latent space. The third stage merges aligned text representations with non-textual features for training a new ID-based model. The final two stages are executed alternately, as denoted by the green dotted line.
@@ 4.1. DAP: Domain-Adaptive Pre-training
4.1.1. Language Model Pre-training
To enhance the PLM’s understanding of dataset-specific texts, we construct a corpus from item-side texts, rich in detailed explanations and descriptions that go beyond mere hard prompts and categorical features. Leveraging this corpus for further pre-training, we utilize the Masked Language Model Loss (MLM) (Devlin et al., 2018), which masks a portion of tokens within a sentence and calculates the loss for each masked token prediction. For instance, the input text ”This ballet skirt is nice for dress up play for ballerinas…,” may be masked as ”This ballet skirt is [MASK] for dress-up play for [MASK]…”, forcing the PLM to predict the mask tokens and give the MLM loss as:
@@ 4.1.1. Language Model Pre-training
Simultaneously, to enhance item text representations and mitigate the anisotropy issue (Ethayarajh, 2019; Li et al., 2020)—where learned representations are constrained to a narrow cone in vector space, reducing expressiveness—we incorporate SimCSE (Gao et al., 2021) into our framework. SimCSE, an unsupervised technique, employs contrastive learning, treating a text as its own positive pair by feeding it into the PLM twice with standard dropout as the only noise. By drawing positive pairs closer, the PLM is compelled to discern the meanings of texts amidst noise, thereby improving its comprehensive capacity. Moreover, distancing the representations of distinct texts promotes a more uniform distribution of representations in vector space, addressing the anisotropy and yielding superior text representations. This process can be formulated as:
(5) $\small\mathcal{L}_{SimCSE}=-\sum_{i\in\mathcal{I}}\log\frac{\exp{\left(sim_{1}\left(\boldsymbol{h}_{i},\boldsymbol{h}^{\prime}_{i}\right)/\tau_{1}\right)}}{\sum_{j\in\mathcal{I}}\exp\left(sim_{1}\left(\boldsymbol{h}_{i},\boldsymbol{h}^{\prime}_{j}\right)/\tau_{1}\right)},$
where $sim_{1}(\cdot)$ denotes the cosine similarity, $\tau_{1}$ denotes the temperature that scales the similarity measure, and $\boldsymbol{h}_{i}$ and $\boldsymbol{h}^{\prime}_{i}$ denote the representations of item $i$’s text.
@@
4.1.2. ID-Based Model Pre-training
To provide a reference for steering the capabilities of language models towards recommendation tasks, we develop an ID-based model that processes exclusively categorical or numerical features. The model’s prediction function is formulated as:

### v3

@@ Abstract.
Click-Through Rate (CTR) prediction holds a paramount position in recommender systems. The prevailing ID-based paradigm underperforms in cold-start scenarios due to the skewed distribution of feature frequency. Additionally, the utilization of a single modality fails to exploit the knowledge contained within textual features. Recent efforts have sought to mitigate these challenges by integrating Pre-trained Language Models (PLMs). They design hard prompts to structure raw features into text for each interaction and then apply PLMs for text processing. With external knowledge and reasoning capabilities, PLMs extract valuable information even in cases of sparse interactions. Nevertheless, compared to ID-based models, pure text modeling degrades the efficacy of collaborative filtering, as well as feature scalability and efficiency during both training and inference. To address these issues, we propose Cost-Efficient Language Model Alignment (CELA) for CTR prediction. CELA incorporates item textual features and language models while preserving the collaborative filtering capabilities of ID-based models. This model-agnostic framework can be equipped with plug-and-play textual features, with item-level alignment enhancing the utilization of external information while maintaining training and inference efficiency. Through extensive offline experiments, CELA demonstrates superior performance compared to state-of-the-art methods. Furthermore, an online A/B test conducted on an industrial advertising recommender system showcases its practical effectiveness, solidifying the potential for real-world applications of CELA. Codes are available at https://github.com/pepsi2222/CELA.
@@ 1. Introduction
Figure 1. ID-Based Models vs. Pre-trained Language Models (PLMs) for CTR prediction. The former converts tabular data into one-hot vectors, capturing collaborative signals without external knowledge. The latter populates data into a hard prompt, and then feeds truncated texts into the PLM, leveraging external world knowledge for data interpretation.
The ID-based paradigm has garnered considerable success. Notable examples include DeepFM (Guo et al., 2017), PNN (Qu et al., 2016), and DCN (Wang et al., 2017), which focus on high-order explicit feature crossing, while DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) are geared towards mining behavior sequences. This paradigm, illustrated on the left side of Figure 1, entails encoding user and item features into sparse one-hot vectors, subsequently transforming into dense embedding vectors via a lookup process. Latent co-occurrences among features are discerned via a feature interaction layer. Over the past decade, the advancement of these models has centered on developing more sophisticated and scenario-adapted feature interaction layers as a means to enhance overall performance. Despite their prevalence, ID-based models exhibit certain limitations: 1) Dependence on historical data: This dependence can result in suboptimal modeling in scenarios such as cold-start or long-tail elements, where historical interactions for items or features are limited or absent, leading to inadequate learning of embeddings. 2) Limited feature scope: Primarily focusing on categorical and numerical features, ID-based models overlook other modalities like text, which may harbor untapped knowledge that could potentially enhance the model’s effectiveness.
To alleviate these limitations, works such as PTab (Liu et al., 2022) and CTRL (Li et al., 2023) have integrated Pre-trained Language Models (PLMs) into recommender systems. These models employ hard prompts to structure raw features for each interaction, subsequently processed through PLMs for recommendation purposes, as shown on the right side of Figure 1. By adopting natural language to represent features instead of one-hot vectors, this approach preserves the semantic integrity of features. This enables PLMs to apply their external knowledge and reasoning capabilities, effectively capturing valuable information even in cases of sparse interactions. For example, ID-based models struggle to learn embeddings for infrequently interacted items like bread, whereas PLMs utilize external knowledge to classify rice as a type of food, akin to toast. Concurrently, this natural language representation facilitates the integration of textual features, such as item descriptions, thereby augmenting the recommender system’s effectiveness.
However, these text-enhanced models face inherent limitations that hinder their practical deployment, shown in Figure 1. First, modeling based exclusively on pure text may not be ideally suited for capturing collaborative signals, which often results in reduced performance (Li et al., 2023). Second, scalability issues emerge as the number of features increases, potentially leading to constructed text exceeding the language model’s maximum token limit and thus losing valuable information, particularly when the sequence of user behaviors is extensively long. Furthermore, it becomes impractical to delineate user behaviors in a detailed manner, including the specification of item features for each behavior (Yang et al., 2024). Last but not least, to bridge the gap between PLM outputs and recommendation tasks, PLMs are usually fine-tuned at the interaction-level (Liu et al., 2022; Geng et al., 2022; Li et al., 2023; Muhamed et al., 2021), yet this process is notably time-intensive. Such inefficiency inherent in PLM text processing extends to inference as well (Liu et al., 2022; Geng et al., 2022; Lin et al., 2024; Bao et al., 2023).
@@
To overcome these challenges, we propose the Cost-Efficient Language Model Alignment (CELA) for CTR prediction. To mitigate cold-start issues, CELA expands its feature scope by leveraging item text features and PLMs to generate more accurate item and user profiles. To maintain optimality and reduce inference latency, CELA integrates ID-based models, sustaining the efficacy and efficiency of collaborative filtering. For scalability, CELA shifts from the hard prompt mode to using PLMs solely as encoders for item textual features. Furthermore, to reduce training costs, CELA fine-tunes PLMs through item-level alignment. CELA is structured as a three-phase paradigm: 1) Domain-Adaptive Pre-training stage, the PLM is further pre-trained on a corpus of item textual features, tailoring it to the specific dataset; 2) Recommendation-Oriented Modal Alignment stage, contrastive learning is employed to align the PLM with the item ID embedding table of a developed ID-based model. This alignment is at the item level, ensuring minimal training overhead; and 3) Multi-Modal Feature Fusion stage, aligned text representations of each item are cached, facilitating efficient access via item identifiers, thereby reducing inference costs. Then these text representations are integrated into a new ID-based model, along with non-textual features, enabling their interplay to identify potential click patterns. The final two stages are structured for alternate execution, ensuring ongoing alignment of the PLM with the progressively refined embedding table.
@@
  • We propose a novel paradigm for integrating item textual features with ID-based models. This approach is model-agnostic and necessitates minimal modifications to the existing network architecture, rendering it conveniently applicable.
@@
  • Comprehensive offline experiments on public and industrial datasets demonstrate the effectiveness of CELA. An online A/B test within an industrial advertising recommender system further validates CELA’s real-world efficacy and feasibility.
@@ 2.1. ID-Based Models for CTR Prediction
ID-based models have been predominant in industry recommendations over the past decade. They capture the characteristics of the items and users by one-hot vectors and learn the collaborative signals to predict the likelihood of a user clicking on an item (Lian et al., 2020). Their architectures are meticulously crafted, evolving into increasingly intricate ones. FM (Rendle, 2010) stands out as one of the earliest machine learning models to incorporate second-order feature interactions. Both WideDeep (Cheng et al., 2016) and DeepFM (Guo et al., 2017) incorporate deep components to learn high-level feature interactions. PNN (Qu et al., 2016) extends the feature interaction function and tightly couples the interaction of first-order and second-order features. AutoInt (Song et al., 2019) and DESTINE (Xu et al., 2021) incorporate self-attention mechanisms to weight different feature interactions. DCN (Wang et al., 2017), DCNv2 (Wang et al., 2021), and EDCN (Chen et al., 2021) capture meaningful feature crossings automatically, eliminating the need for manual feature engineering or exhaustive searching. DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) model user behavior sequences, aiming to capture the relationships between the target item and historical behaviors.
@@ 2.2. Text-Enhanced Models for CTR Prediction
Language models have made remarkable strides, achieving unprecedented natural language understanding and generation. Pre-trained on extensive datasets and utilizing sophisticated architectures like Transformer (Vaswani et al., 2017), these models support diverse applications such as chatbots (Du et al., 2022), translation (Zhang et al., 2022), and content creation (Chung et al., 2022). Highlighting their potential in semantic comprehension, language models are considered promising for recommendation. Analyzing extra-textual elements such as product descriptions, these models may have the capability to discern user intent and item nuances, thereby enhancing ID-based ones. P5 (Geng et al., 2022) redefines tasks as natural language processing (NLP) tasks, structuring interactions and metadata in natural language and inputting them into T5 (Raffel et al., 2020) to generate target responses. CTR-BERT (Muhamed et al., 2021) concatenates all features with their names, undergoes pre-training and distillation with BERT (Devlin et al., 2019), and encodes user and item text to generate representations for a late fusion MLP. PTab (Liu et al., 2022) is pre-trained on BERT using a ”field: value” corpus from tabular data, then fine-tuned for the CTR task with a classifier head. UniSRec (Hou et al., 2022) employs a frozen BERT as a text encoder and utilizes a Transformer architecture to model user behaviors. CTRL (Li et al., 2023) uses contrastive learning to align representations between collaborative and semantic models at the interaction level. The collaborative model, now enriched with semantic information, is fine-tuned and can be deployed independently.
@@ 3.2. Embedding Layer
To compress those sparse, high-dimensional vectors into dense, lower-dimensional ones, the embedding layer is utilized. It acts as a lookup table, denoted by $\mathbf{E}^{(\cdot)}\in\mathbb{R}^{C^{(\cdot)}\times d}$ , which maps each categorical feature into a dense vector of a fixed size $d$. Here, $d$ represents the embedding size. Let $\boldsymbol{x}^{u}$ denote user profile, $\boldsymbol{x}^{h}$ denote browsing history, $\boldsymbol{x}^{c}$ denote contextual information, $\boldsymbol{x}^{i,n}$ denote non-textual features of candidate item $i$ and $\boldsymbol{x}^{h,n}$ denote non-textual features in user history. This process can be formalized as:
@@ 4. Methodology
Figure 2. The overall framework of CELA. The first stage pre-trains a PLM on domain-specific item texts. In the second stage, an ID-based model is developed and item text representations from the PLM are aligned with the item-side feature embeddings of the ID-based model in latent space. The third stage merges aligned text representations with non-textual features for training a new ID-based model. The final two stages are executed alternately, as denoted by the green dotted line.
@@
  • Domain-Adapted Pre-training (DAP). It involves further pre-training of a PLM on a corpus comprising item-side text specific to the recommendation dataset, to better adapt the PLM to the unique characteristics of the dataset.
@@
Notably, the last two stages can be executed alternatively, where the PLM is continually aligned to the more advanced ID embedding table.
@@ 4.1. DAP: Domain-Adaptive Pre-training
To enhance the PLM’s understanding of dataset-specific texts, we construct a corpus from item-side texts, rich in detailed item descriptions, as illustrated in Appendix A. Leveraging this corpus for further pre-training, we utilize the Masked Language Model Loss (MLM) (Devlin et al., 2019), which masks a portion of tokens within a sentence and calculates the loss for each masked token prediction. For instance, the input text ”This ballet skirt is nice for dress-up play for ballerinas…,” may be masked as ”This ballet skirt is [MASK] for dress-up play for [MASK]…”, forcing the PLM to predict the mask tokens and give the MLM loss as:
@@
Simultaneously, to enhance item text representations and mitigate the anisotropy issue (Ethayarajh, 2019; Li et al., 2020)—where learned representations are constrained to a narrow cone in vector space, reducing expressiveness—we incorporate SimCSE (Gao et al., 2021) into our framework. SimCSE, an unsupervised technique, employs contrastive learning, treating a text as its positive pair by feeding it into the PLM twice with standard dropout as the only noise. By drawing positive pairs closer, the PLM is compelled to discern the meanings of texts amidst noise, thereby improving its comprehensive capacity. Moreover, distancing the representations of distinct texts promotes a more uniform distribution of representations in vector space, addressing the anisotropy and yielding superior text representations. This process can be formulated as:
(5) $\small\mathcal{L}_{SimCSE}=-\sum_{i\in\mathcal{I}}\log\frac{\exp{\left(sim\left(\boldsymbol{h}_{i},\boldsymbol{h}^{\prime}_{i}\right)/\tau_{1}\right)}}{\sum_{j\in\mathcal{I}}\exp\left(sim\left(\boldsymbol{h}_{i},\boldsymbol{h}^{\prime}_{j}\right)/\tau_{1}\right)},$
where $sim(\cdot)$ denotes the similarity function, $\tau_{1}$ denotes the temperature scaling the similarity measure, and $\boldsymbol{h}_{i}$ and $\boldsymbol{h}^{\prime}_{i}$ denote the representations of item $i$’s text.
@@
4.2. ROMA: Recommendation-Oriented Modal Alignment
Despite the profound domain understanding that the PLM achieves through pretraining, a significant gap persists between their output space and that of collaborative filtering. This discrepancy arises because the capabilities of the PLM extend well beyond what recommendation tasks require. Without targeted training and constraints, the outputs of the PLM may not align with the needs of recommender systems. Consequently, it is crucial to refine and orient the PLM’s outputs towards recommendation tasks, a process referred to as alignment. This essential alignment, which aims to provide a representation of the item’s textual feature that can be seamlessly integrated with other features for the next stage, is illustrated in the central portion of Figure 2. The substantial number of parameters in the PLM and the extensive volume of interactions in the dataset render interaction-level alignment inefficient (Li et al., 2023; Liu et al., 2022; Geng et al., 2022). Consequently, we propose item-level alignment as a solution.
4.2.1. ID-Based Model
To provide a reference for steering capabilities of the PLM towards recommendation tasks, we develop an ID-based model that processes exclusively categorical or numerical features. The model’s prediction function is formulated as: